# LoRA Meets Dropout under a Unified Framework

### v1

@@ Abstract
Sheng ${}^{\heartsuit}$thanks:   Equal Contribution. , Liheng ${}^{\heartsuit\ast}$, Jiyue ${}^{\spadesuit}$, Boyang ${}^{\spadesuit}$, Lingpeng ${}^{\heartsuit}$, Chuan ${}^{\heartsuit}$$^{\heartsuit}$ The University of Hong Kong, ${}^{\spadesuit}$ The Chinese University of Hong Kong {u3009618, clh648}@connect.hku.hk, jiangjy@link.cuhk.edu.hk, byxue@se.cuhk.edu.hk, {lpk, cwu}@cs.hku.hk
@@ 3.2 A Unified Framework
For better generalization, a robust model needs to learn noise-resilient features. Hence, dropping position, determining where to inject noise, emerges as a primary consideration in designing dropout methods. For example, dropping inputs acts like data augmentation, dropping outputs encourages an ensemble of sub-classifiers, and dropping intermediate representations disrupts the co-adaptation of neighboring neurons. For a transformer layer depicted in Figure 1, DropKey, DropAttention and HiddenCut respectively drop attention logits, weights and hidden representations, covering the self-attention mechanism and feed-forward module. Additionally, the same dropping position may perform differently in full finetuning and LoRA scenarios. In full finetuning, weights directly associated with the dropping position are continuously adjusted for better noise resilience, while this adaptation is more implicit for LoRA. Specifically, LoRA, typically applied to the key and value projection matrices (Hu et al., 2021), requires multiple intermediate calculations to influence attention logits and weights, while even requires inter-module computation for hidden representations. This disparity may potentially affect the effectiveness of existing dropout methods in LoRA scenarios. Notably, distinct dropping positions do not necessarily indicate differences. In specific cases, different positions may also exhibit similar features, as discussed in Sec. 3.1.
@@ 4.4 Superiority on More NLU and NLG Tasks
We further generalize HiddenKey to two extra NLU datasets, namely CoLA and QNLI. As shown in Table 4, HiddenKey steadily achieves 1.95 and 0.81 performance improvement over baselines on both of the datasets. These reconfirm the superior performance of HiddenKey in NLU tasks.

### v2

@@ Abstract
Sheng Wangthanks:   Equal Contribution. , Liheng Chen♡∗, Jiyue Jiang, Boyang Xue, Lingpeng Kong, Chuan Wu The University of Hong Kong, The Chinese University of Hong Kong {u3009618, clh648}@connect.hku.hk, jiangjy@link.cuhk.edu.hk, byxue@se.cuhk.edu.hk, {lpk, cwu}@cs.hku.hk
@@ 3.2 A Unified Framework
For better generalization, a robust model needs to learn noise-resilient features. Hence, dropping position, determining where to inject noise, emerges as a primary consideration in designing dropout methods. For example, dropping inputs acts like data augmentation, dropping outputs encourages an ensemble of sub-classifiers, and dropping intermediate representations disrupts the co-adaptation of neighboring neurons. For a transformer layer depicted in Figure 1, DropKey, DropAttention and HiddenCut respectively drop attention logits, weights and hidden representations, covering the self-attention mechanism and feed-forward module. Additionally, the same dropping position may perform differently in full finetuning and LoRA scenarios. In full finetuning, weights located in the dropping position are directly adjusted for better noise resilience. However, this adaptation is more implicit for LoRA, because the directly associated weights with the dropping position are frozen. Specifically, LoRA, typically applied to the key and value projection matrices (Hu et al., 2021), requires multiple intermediate calculations (e.g., softmax) to influence attention logits and weights (i.e., the dropping positions for DropAttention and DropKey), while even requires inter-module computation for hidden representations. This disparity may potentially affect the effectiveness of existing dropout methods in LoRA scenarios. Notably, distinct dropping positions do not necessarily indicate differences. In specific cases, different positions may also exhibit similar features, as discussed in Sec. 3.1.
@@ 4.4 Superiority on More NLU and NLG Tasks
We further generalize HiddenKey to two extra NLU datasets, namely CoLA and QNLI. As shown in Table 4, HiddenKey steadily achieves 1.95 and 0.81 performance improvement over baselines on both of the datasets, reconfirming HiddenKey’s superiority in NLU tasks.
@@ 7 Ethics Statement
8 Acknowledgement
This work was supported in part by Hong Kong Innovation and Technology Support Programme Platform Research Project fund (ITS/269/22FP), the joint research scheme of the National Natural Science Foundation of China (NSFC) and Hong Kong Research Grants Council (RGC) (under grant N_HKU714/21), and RGC grants 17204423 and C7004-22G (CRF).
@@ Appendix F Finetuning dynamics
Appendix G Statistical Significance Test
Model (Benchmark) Method p-value
Baseline 0.080
RoBERTa DropKey 0.059
(GLUE) HiddenCut 0.024
HiddenKey- 0.031
Baseline 0.044
GPT-2 DropKey 0.052
(E2E) HiddenCut 0.041
HiddenKey- 0.043
Table 8: P-values of HiddenKey versus alternative methods.
To assess the statistical significance of the results presented in Table 1 and Table 2, we calculate the p-values of HiddenKey comparing against alternative approaches, averaged on the benchmarks. As shown in Table 8, the obtained p-values, all below 0.1 with the majority falling below 0.05, strongly indicate the statistical significance of HiddenKey’s superiority.